OpenStudy (anonymous):

3. Solve the system of equations using linear combination. a + c = 9 and 8a +4.5c = 58

OpenStudy (anonymous):

a + c = 9 c = 9-a 8a + 4.5(9-a) = 58 8a + 40.5 -4.5a = 58 3.5a = 17.5 a = 5 5 + c = 9 c= 4 a= 5 c= 4